OpenStudy (anonymous):

suppose you need to hang a 300 +/-10 g on the weight of the hanger to observe a standing wave. what are the values of T and standard deviation of T?

OpenStudy (anonymous):

hey eashmore.. for this problem i used T=mg for each value . but im unsure about the standard deviation of T

OpenStudy (anonymous):

Are you sure that T is tension? The question mentions a standing wave, which makes me think of simple harmonic motion. T could be the period of oscillation.

OpenStudy (anonymous):

i think u r right buit i do not know what equation only reguires mass ( only thing we r given)

OpenStudy (anonymous):

Let's assume it is tension. I'm 90% sure the following is correct. Be careful. To determine the standard deviation, we need to use the following formula\[\sigma(T) = \sqrt{ \left({\partial T \over \partial g} \sigma(g)\right)^2 + \left( {\partial T \over \partial m} \sigma(m) \right)^2 }\]The uncertainty of gravity can be said to be zero, therefore \(\sigma(g) = 0\). The uncertainty of mass \(\sigma(m)\) is 10/300 (We want it as a percentage.) Therefore, the standard deviation of T is\[\sigma(T) = \sqrt{ \left( g \cdot \left(10 \over 300 \right) \right)^2}\]

OpenStudy (anonymous):

in the denominator is that a 2m

OpenStudy (anonymous):

i dont get ur question well but as to my understanding,it is a standing wave meaning there is no propagation of energy so there arent any vibrations created in the string so mg=T as there is acc or force and hence energy movement involved here